# On graphs with large third eigenvalue

### v3

@@ 3.6 Pivalous graphs
First, $\lambda_{1}=n/2+1$ represents the regularity. Otherwise, we have $\lambda=(\omega^{t})^{-n/4}\frac{(\omega^{t})^{n/2+1}-1}{\omega^{t}-1}$. We now consider cases on $t\pmod{4}$. {outline} \1[0:] $\lambda=\frac{\omega^{t}-1}{\omega^{t}-1}=+1$. \1[1:] $\lambda=-i\frac{-\omega^{t}-1}{\omega^{t}-1}=i\frac{\omega^{t}+1}{\omega^{t}-1}=+\cot(\frac{t\pi}{n})$ by Lemma 3.7. \1[2:] $\lambda=-\frac{\omega^{t}-1}{\omega^{t}-1}=-1$. \1[3:] $\lambda=i\frac{-\omega^{t}-1}{\omega^{t}-1}=-\cot(\frac{t\pi}{n})$. Note that we get $+\cot(\frac{\pi}{n})=-\cot(\frac{(n-1)\pi}{n})$. These are the second and third eigenvalues by size comparison to the rest, hence $\lim_{n\to\infty}\frac{\lambda_{3}(Pi_{n})}{n}\to\frac{1}{\pi}$. ∎
@@
By observing their graphon representations in the sense of Lovász and Szegedy [9], we can verify that these graphs converge in cut metric to a graphon $W$ where $W(x,y)=1$ if $0\leq x-y\leq\frac{1}{4}\pmod{1}$ or $0\leq y-x\leq\frac{1}{4}\pmod{1}$. Borgs et al. [3] proved the limit of the spectrum is the spectrum of the limit, hence we can derive the spectrum of $W$. Alternatively, note that the eigenfunctions are $f_{k}(x)=\exp(2i\pi kx)$ for $k\in\mathbb{Z}$.
@@ 4 Extremising $\lambda_{n-1}$ over weighted graphs
In other words, we consider the extension of the problem to weighted undirected graphs (possibly with loops) with weights in $[0,1]$. Remark that the function $\lambda_{k}:\mathcal{S}_{n}\to\mathbb{R}$ is well defined and if we equip $\mathcal{S}_{n}$ with the metric induced by the operator norm $||\cdot||_{\text{opp}}$ on $n\times n$ matrices, Weyl’s inequality gives:
$|\lambda_{k}(M)-\lambda_{k}(N)|\leq||M-N||_{\text{opp}}.$
@@
$-\frac{n}{3}\leq\inf_{M\in\mathcal{S}_{n}}\lambda_{n-1}(M).$
In this section, we make progress to answering Question 4.1 for $k=n-1$ and motivate Conjecture 4.2. Firstly, we remark the following lemma, similar to that of Theorem 1.2 of Csikvári [4].
@@ 4.1 Implications of the existence of an extremiser of $\lambda_{n-1}$ with $\lambda_{n}$ simple
where $O_{m}$ denotes the $m\times m$ matrix of zeros, $J_{m,l}$ denotes the $m\times l$ matrix of ones and $X,Y,W$ are $P\times Z$, $N\times Z$ and $Z\times Z$ matrices with $[0,1]$ entries respectively, with $\frac{1}{\sqrt{P}}X^{T}\mathbb{j}_{P}=\frac{1}{\sqrt{N}}Y^{T}\mathbb{j}_{N}$. If $\mathbb{w}$ denotes the eigenvector of $M$ corresponding to $\lambda_{n-1}$, then $\mathbb{w}^{T}=\begin{pmatrix}\frac{\mathbb{j}_{P}^{T}}{\sqrt{2P}}&-\frac{\mathbb{j}_{N}^{T}}{\sqrt{2N}}&\mathbb{0}_{Z}^{T}\end{pmatrix}$ and $\lambda_{n-1}=-\sqrt{PN}$.
@@
Suppose that $M$ is a simple minimiser with eigenvalues $\lambda_{n}(M)<\lambda_{n-1}(M)$ and corresponding orthonormal eigenvectors $\mathbb{v,w}$. Suppose that $E$ is an $n\times n$ matrix such that $\exists\eta>0$ with $M+\varepsilon E\in\mathcal{S}_{n}$ we have $\forall 0<\varepsilon<\eta$. Then, by the min-max principle:
$\lambda_{n-1}(M+\varepsilon E)-\lambda_{n-1}(M)$ $\leq\max_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v,w}}\\ ||\mathbb{x}||=1\end{subarray}}\left(\mathbb{x}^{T}(M+\varepsilon E)\mathbb{x}\right)-\lambda_{n-1}(M)$
@@
$\lambda_{n-1}(M+\varepsilon E)-\lambda_{n-1}(M)$ $\leq\max_{\alpha\in[-1,1]}\left(\alpha^{2}(\lambda_{n}(M)-\lambda_{n-1}(M)+\varepsilon(S-R))+2\varepsilon\alpha Q+\varepsilon R\right)$
@@
Note that if $R<0$ then (*) is negative for all $\varepsilon>0$ sufficiently small. Thus, if there exists such an $E$ with $M+\varepsilon E\in\mathcal{S}_{n}$ for $\varepsilon$ sufficiently small and $\mathbb{w}^{T}E\mathbb{w}<0$, then $\lambda_{n-1}(M+\varepsilon E)<\lambda_{n-1}(M)$, contradicting the minimality of $\lambda_{n-1}(M)$ amongst elements of ${\mathcal{S}}_{n}$.
@@
For such $E$, $R=\mathbb{w}^{T}E\mathbb{w}=-2w_{i}w_{j}<0$, so by the above, $M$ is not minimal (yielding a contradiction). Hence, if $i\neq j$ are such that $(i,j)\in W_{+}\times W_{+}\cup W_{-}\times W_{-}$, then $M_{ij}=0$.
@@
we again have $R=\mathbb{w}^{T}E\mathbb{w}=2w_{i}w_{j}<0$ giving a contradiction. Thus, if $i,j$ are such that $(i,j)\in W_{+}\times W_{-}\cup W_{-}\times W_{+}$, then $M_{ij}=1$. Hence, $M$ must have form:
@@
If $M$ is a simple minimiser and some $M^{\prime}\in\mathcal{S}_{n}$ has $\lambda_{n}(M^{\prime})\leq\lambda_{n}(M)$ as well as $\lambda_{n-1}(M^{\prime})\leq\lambda_{n-1}(M)$, then $M^{\prime}$ is also a simple minimiser since $\lambda_{n}(M^{\prime})\leq\lambda_{n}(M)<\lambda_{n-1}(M)=\lambda_{n-1}(M^{\prime})$. In the following, we assume the existence of such an $M$ and repeatedly construct another simple minimiser $M^{\prime}$ from $M$ with nicer properties.
@@
where $C$ is a $(Z_{1}+Z_{2})\times(Z_{1}+Z_{2})$ matrix such that $M+E\in\mathcal{S}_{n}$. Remark that $(M+E)\mathbb{w}=\lambda_{n-1}(M)\mathbb{w}$, so if $\lambda_{n}(M+E)\leq\lambda_{n}(M)$ then $M+E$ is also a simple minimiser. We may bound:
@@
Remark that $(M+E)\mathbb{w}=-\sqrt{PN}\mathbb{w}$. Moreover, we have that:
@@
If this is non-positive, then applying this operation to each row which is equal to the $j$-th row of $\begin{pmatrix}X_{1}^{T}&Y_{1}^{T}\end{pmatrix}$ gives another simple minimiser and strictly reduces the total number of distinct rows (contradicting maximality). Otherwise, this quantity is negative and so switching the roles of $i$ and $j$ in the above (i.e. copying the row $j$ into row $i$ and column $j$ into row $i$) yields:
@@
So, picking $\varepsilon=\min\{1-X_{ip},X_{iq}\}$, we have strictly increased the number of entries in $\{0,1\}$, contradicting maximality. Thus, we may suppose that there is at most one entry of each of $X_{1}^{T},Y_{1}^{T}$ which is not 00 nor $1$.
Of these matrices, further pick any $M$ with the longest combined length of: the runs of 1 entries which begin at start of each (identical) row of $X_{1}^{T}$ and $Y_{1}^{T}$ as well as the runs of 0 entries which finish at the end of each row of $X_{1}^{T}$ and $Y_{1}^{T}$. For example, the following gives $1+3+2+0=6$.
@@
Suppose that for some $p<q\leq P$ we have $(X_{1})_{ip}<1$ and $(X_{1})_{iq}=1$. Select the first $p$ for which this occurs. Then as above, we have that $\lambda_{n}(M+(1-(X_{1})_{ip})E)-\lambda_{n}(M)=\sum_{i=1}^{Z_{1}}2(1-(X_{1})_{ip})(\mathbb{v}_{Z_{1}})_{i}((\mathbb{v}_{P})_{p}-(\mathbb{v}_{P})_{q})\leq 0$. The last eigenvalue does not increase, whilst the number of entries in $\{0,1\}$ stays constant and the length of the run of 1’s at the start of each row of $X_{1}^{T}$ strictly increases, which gives a contradiction. Remark that this does not affect any other starting run of 1 or ending run of 0, since if $(X_{1})_{ip}=0$ then having $(X_{1})_{iq}=1$ after blocks this from connecting to the end of $X_{1}^{T}$. Thus, we have that all the 1 entries of $X_{1}^{T}$ lie in the initial run. Applying a similar argument to the run of 0 at the end of each row of $X_{1}^{T}$ and repeating for $Y_{1}^{T}$, we may conclude that each row of $\begin{pmatrix}X_{1}^{T}&Y_{1}^{T}\end{pmatrix}$ has form:
@@
The matrix $\tilde{M}$ is given by a vertex multiplication of the following $11\times 1$ 11 matrix:
@@
Remark that the complement of $H_{a,b}$ is isomorphic to the vertex multiplication of $T$ by (for example) $[a,0,0,b,a,0,0,b,a,b,0]$. The non-zero eigenvalues of the general vertex multiplication of $T$ are given by the roots of a sextic polynomial, which we have included in the Appendix. In light of Lemma 4.3, we conjecture that:
@@ 4.2 Extremisers of $\lambda_{n-1}$ with $\lambda_{n}$ repeated
$0\leq\lambda_{n-1}(M+E)-\lambda_{n-1}(M)$ $\leq\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v},\mathbb{w}}\\ ||\mathbb{x}||=1\end{subarray}}(\mathbb{x}^{T}(M+E)\mathbb{x})-\lambda_{n-1}(M)$
$=\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v},\mathbb{w}}\\ ||\mathbb{x}||=1\end{subarray}}(\mathbb{x}^{T}E\mathbb{x})\leq 0.$
@@
Thus if $n\geq 15$, we have $\lambda_{n-1}(M+E)\geq-\frac{n}{2\sqrt{3}}>-\frac{n-2}{3}$. But, the complements of $H_{a,b}$ have $\lambda_{n-1}(\overline{H_{a,b}})=-\frac{|V(\overline{H_{a,b}})|}{3}$, contradict the extremality of $\lambda_{n-1}(M+E)$. The authors have also verified for $5\leq n\leq 14$, there exists weighted graphs with penultimate eigenvalue lower than $-\frac{n}{2\sqrt{3}}$.
@@
$\lambda_{n-1}(M+E)=\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v},\mathbb{w}}\\ ||\mathbb{x}||=1\end{subarray}}\mathbb{x}^{T}(M+E)\mathbb{x}.$ (*)
Let $\tilde{\mathbb{v}}_{n},\tilde{\mathbb{v}}_{n-1},\dots,\tilde{\mathbb{v}}_{1}$ be orthonormal eigenvectors of $M+E$, with corresponding eigenvalues $\tilde{\lambda}_{n}=\tilde{\lambda}_{n-1},\dots\tilde{\lambda}_{1}$. Then, consider the vector space $V=\braket{\mathbb{v,w}}\cap\braket{\mathbb{\tilde{v}}_{1},\dots\mathbb{\tilde{v}}_{n-1}}$. Note that $V$ has dimension at least 1, hence there exists some $\mathbb{z}\in V$ with $||\mathbb{z}||=1$. Thus, if $\mathbb{x}=\sum_{i=1}^{n-1}a_{i}\mathbb{\tilde{v}}_{i}$, then:
@@
Remark that:
$\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v,w}}\\ ||\mathbb{x}||=1\end{subarray}}\mathbb{x}^{T}E\mathbb{x}=\sup_{\begin{subarray}{c}\alpha,\beta\\ \alpha^{2}+\beta^{2}=1\end{subarray}}(S\alpha^{2}+2Q\alpha\beta+R\beta^{2}).$
@@
$\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v,w}}\\ ||\mathbb{x}||=1\end{subarray}}\mathbb{x}^{T}E\mathbb{x}=\sup_{\begin{subarray}{c}\alpha,\beta\\ \alpha^{2}+\beta^{2}=1\end{subarray}}\left(S\left(\alpha+\frac{Q}{S}\beta\right)^{2}+\left(\frac{RS-Q^{2}}{S}\right)\beta^{2}\right),$
which is less than or equal to 00 if and only if $S<0$ and $RS\geq Q^{2}$ (implying that $R\leq 0$). Identical logic applies when $R\neq 0$.
@@
$\sup_{\begin{subarray}{c}\mathbb{x}\in\braket{\mathbb{v,w}}\\ ||\mathbb{x}||=1\end{subarray}}\mathbb{x}^{T}E\mathbb{x}=\sup_{\begin{subarray}{c}\alpha,\beta\\ \alpha^{2}+\beta^{2}=1\end{subarray}}(2Q\alpha\beta)=|Q|,$
@@
$E_{xy}=\begin{cases}-M_{ii}&x=y\\ 0&\text{otherwise.}\end{cases}$
@@
There exists a minimiser $M$ of $\lambda_{n-1}$ with at most 1 entry of each row and column not 00 nor $1$.
@@
Let $M$ be a minimiser of $\lambda_{n-1}$ with diagonal entries all zero and the least number of entries that are not 00 nor $1$. Suppose by way of contradiction some row (or equivalently column) of $M$ has at least 2 entries $M_{ab},M_{ac}\in(0,1)$ for some $a,b,c$ distinct. Consider symmetric matrices $E$ with 0 entries everywhere other than at $E_{ab},E_{ac},E_{bc}$ (and $E_{ba}$, $E_{ca}$, $E_{cb}$). Then, we have that:
@@
Regardless of the existence of a simple minimiser, there exists a matrix $N$ with $0,1$ entries such that for all $M\in\mathcal{S}_{n}$:

### v4

@@ 3.6 Pivalous graphs
First, $\lambda_{1}=n/2+1$ represents the regularity. Otherwise, we have $\lambda=(\omega^{t})^{-n/4}\frac{(\omega^{t})^{n/2+1}-1}{\omega^{t}-1}$. We now consider cases on $t\pmod{4}$.
  • 0:
    $\lambda=\frac{\omega^{t}-1}{\omega^{t}-1}=+1$.
  • 1:
    $\lambda=-i\frac{-\omega^{t}-1}{\omega^{t}-1}=i\frac{\omega^{t}+1}{\omega^{t}-1}=+\cot(\frac{t\pi}{n})$ by Lemma 3.7.
  • 2:
    $\lambda=-\frac{\omega^{t}-1}{\omega^{t}-1}=-1$.
  • 3:
    $\lambda=i\frac{-\omega^{t}-1}{\omega^{t}-1}=-\cot(\frac{t\pi}{n})$.
Note that we get $\cot(\frac{\pi}{n})=-\cot(\frac{(n-1)\pi}{n})$. These are the second and third eigenvalues by size comparison to the rest, hence $\lim_{n\to\infty}\frac{\lambda_{3}(Pi_{n})}{n}\to\frac{1}{\pi}$. ∎
@@
By observing their graphon representations in the sense of Lovász and Szegedy [9], we can verify that these graphs converge in cut metric to a graphon $W$ where $W(x,y)=1$ if $0\leq x-y\leq\frac{1}{4}\pmod{1}$ or $0\leq y-x\leq\frac{1}{4}\pmod{1}$. Borgs et al. [BorgsGraphon2] proved the limit of the spectrum is the spectrum of the limit, hence we can derive the spectrum of $W$. Alternatively, note that the eigenfunctions are $f_{k}(x)=\exp(2i\pi kx)$ for $k\in\mathbb{Z}$.
@@ 4 Extremising $\lambda_{n-1}$ over weighted graphs
In other words, we consider the extension of the problem to weighted undirected graphs (possibly with loops) with weights in $[0,1]$. We remark that the function $\lambda_{k}:\mathcal{S}_{n}\to\mathbb{R}$ is well defined and if we equip $\mathcal{S}_{n}$ with the metric induced by the operator norm $||\cdot||_{\text{opp}}$ on $n\times n$ matrices, Weyl’s inequality gives:
[…]
@@
$\frac{n}{3}\leq\inf_{M\in\mathcal{S}_{n}}\lambda_{n-1}(M).$
In this section, we make progress to answering Question 4.1 for k=n1k=n-1 and motivate Conjecture 4.2. Firstly, we remark the following lemma, similar to that of Theorem 1.2 of Csikvári [4].
@@ 4.1 Implications of the existence of an extremiser of $\lambda_{n-1}$ with $\lambda_{n}$ simple
where $O_{m}$ denotes the $m\times m$ matrix of zeros, $J_{m,l}$ denotes the $m\times l$ matrix of ones and $X,Y,W$ are $P\times Z$, $N\times Z$ and $Z\times Z$ matrices with $[0,1]$ entries respectively, with $\frac{1}{\sqrt{P}}X^{T}\mathbb{j}_{P}=\frac{1}{\sqrt{N}}Y^{T}\mathbb{j}_{N}$. If $\mathbb{w}$ denotes the unit eigenvector of $M$ corresponding to $\lambda_{n-1}$, then $\mathbb{w}^{T}=\begin{pmatrix}\frac{\mathbb{j}_{P}^{T}}{\sqrt{2P}}&-\frac{\mathbb{j}_{N}^{T}}{\sqrt{2N}}&\mathbb{0}_{Z}^{T}\end{pmatrix}$ and $\lambda_{n-1}=-\sqrt{PN}$.
@@
Suppose that $M$ is a simple minimiser with eigenvalues $\lambda_{n}(M)<\lambda_{n-1}(M)$ and corresponding orthonormal eigenvectors $\mathbb{v,w}$. Suppose that $E$ is an $n\times n$ matrix such that $\exists\eta>0$ with $M+\varepsilon E\in\mathcal{S}_{n}$ for all $0<\varepsilon<\eta$. Then, by the min-max principle:
[…]
@@
$\lambda_{n-1}(M+\varepsilon E)-\lambda_{n-1}(M)$ $=\max_{\alpha\in[-1,1]}\left(\alpha^{2}(\lambda_{n}(M)-\lambda_{n-1}(M)+\varepsilon(S-R))+2\varepsilon\alpha\sqrt{1-\alpha^{2}}Q+\varepsilon R\right)$
$\leq\max_{\alpha\in[-1,1]}\left(\alpha^{2}(\lambda_{n}(M)-\lambda_{n-1}(M)+\varepsilon(S-R))+2\varepsilon\alpha Q+\varepsilon R\right)$
@@
Here, we chose $\alpha$ which maximises the above quadratic.
Claim.
If $E$ is a matrix such that $M+\varepsilon E\in\mathcal{S}_{n}$ for all $\varepsilon$ sufficiently small, then $R\geq 0$.
Proof.
If by way of contradiction $R<0$ then (*) is negative for all $\varepsilon>0$ sufficiently small. Thus, for $\varepsilon$ sufficiently small we have $\lambda_{n-1}(M+\varepsilon E)<\lambda_{n-1}(M)$, contradicting the minimality of $\lambda_{n-1}(M)$ amongst elements of ${\mathcal{S}}_{n}$. ∎
@@
For such $E$, $R=\mathbb{w}^{T}E\mathbb{w}=-2w_{i}w_{j}<0$, so by the claim above, $M$ is not minimal (yielding a contradiction). Hence, if $i\neq j$ are such that $(i,j)\in W_{+}\times W_{+}\cup W_{-}\times W_{-}$, then $M_{ij}=0$.
@@
we again have $R=\mathbb{w}^{T}E\mathbb{w}=2w_{i}w_{j}<0$ giving a contradiction. Thus, if $i,j$ are such that $(i,j)\in W_{+}\times W_{-}\cup W_{-}\times W_{+}$, then $M_{ij}=1$. Hence, $M$ must have the form:
@@
If $M$ is a simple minimiser and some $M^{\prime}\in\mathcal{S}_{n}$ has $\lambda_{n}(M^{\prime})\leq\lambda_{n}(M)$ as well as $\lambda_{n-1}(M^{\prime})\leq\lambda_{n-1}(M)$ (so, by the minimality of $M$, $\lambda_{n-1}(M^{\prime})=\lambda_{n-1}(M)$), then $M^{\prime}$ is also a simple minimiser since $\lambda_{n}(M^{\prime})\leq\lambda_{n}(M)<\lambda_{n-1}(M)=\lambda_{n-1}(M^{\prime})$. In the following, we assume the existence of such an $M$ and repeatedly construct another simple minimiser $M^{\prime}$ from $M$ with nicer properties.
@@
where $C$ is a $(Z_{1}+Z_{2})\times(Z_{1}+Z_{2})$ matrix such that $M+E\in\mathcal{S}_{n}$. We remark that $(M+E)\mathbb{w}=\lambda_{n-1}(M)\mathbb{w}$, so if $\lambda_{n}(M+E)\leq\lambda_{n}(M)$ then $M+E$ is also a simple minimiser. We may bound:
@@
We remark that $(M+E)\mathbb{w}=-\sqrt{PN}\mathbb{w}$. Moreover, we have that:
@@
If this is non-positive, then applying this operation to each row which is equal to the $j$-th row of $\begin{pmatrix}X_{1}^{T}&Y_{1}^{T}\end{pmatrix}$ gives another simple minimiser and strictly reduces the total number of distinct rows (contradicting maximality). Otherwise, this quantity is negative and so switching the roles of $i$ and $j$ in the above yields:
@@
So, picking $\varepsilon=\min\{1-(X_{1})_{ip},(X_{1})_{iq}\}$, we have strictly increased the number of entries in $\{0,1\}$, contradicting maximality. Thus, we may suppose that there is at most one entry of each of $X_{1}^{T},Y_{1}^{T}$ which is not 0 nor $1$.
Of these matrices, further pick any $M$ with the longest combined length of: the runs of 1 entries which begin at the start of each (identical) row of $X_{1}^{T}$ and $Y_{1}^{T}$ as well as the runs of 0 entries which finish at the end of each row of $X_{1}^{T}$ and $Y_{1}^{T}$. For example, the following has a combined length of $1+3+2+0=6$.
@@
Suppose that for some $p<q\leq P$ we have $(X_{1})_{ip}<1$ and $(X_{1})_{iq}=1$. Select the first $p$ for which this occurs. Then as above, we have that $\lambda_{n}(M+(1-(X_{1})_{ip})E)-\lambda_{n}(M)=\sum_{i=1}^{Z_{1}}2(1-(X_{1})_{ip})(\mathbb{v}_{Z_{1}})_{i}((\mathbb{v}_{P})_{p}-(\mathbb{v}_{P})_{q})\leq 0$. The last eigenvalue does not increase, whilst the number of entries in $\{0,1\}$ stays constant and the length of the run of 1’s at the start of each row of $X_{1}^{T}$ strictly increases, which gives a contradiction. We remark that this does not affect any other starting run of 1 or ending run of 0, since if $(X_{1})_{ip}=0$ then having $(X_{1})_{iq}=1$ after blocks this from connecting to the end of $X_{1}^{T}$. Thus, we have that all the 1 entries of $X_{1}^{T}$ lie in the initial run. Applying a similar argument to the run of 0 at the end of each row of $X_{1}^{T}$ and repeating for $Y_{1}^{T}$, we may conclude that each row of $\begin{pmatrix}X_{1}^{T}&Y_{1}^{T}\end{pmatrix}$ has the form:
@@
Thus, the penultimate eigenvalue of $\tilde{M}$ differs from that of a simple minimiser by at most $2\sqrt{n}$. The advantage of working with $\tilde{M}$ is that it has a particularly simple symmetric block form, with entries in $\{0,1\}$. We remark that there are two possibilities for the overlaps between the blocks of ones and zeros of $X_{1}^{T}$ and $X_{2}^{T}$. Either the blocks of ones overlap (i.e. there exists some column index where both $X_{1}^{T}$ and $X_{2}^{T}$ have all entries given by ones) or they do not.
Accounting for both possibilities for the overlaps for $X_{1}^{T}$ and $X_{2}^{T}$ and both possibilities for the overlaps for $Y_{1}^{T}$ and $Y_{2}^{T}$, the matrix $\tilde{M}$ is thus given by a vertex multiplication (possibly with some vertices multiplied by zero) of the following $11\times 11$ matrix:
@@
We remark that the complement of $H_{a,b}$ is isomorphic to the vertex multiplication of $T$ by (for example) $[a,0,0,b,a,0,0,b,a,b,0]$. The non-zero eigenvalues of the general vertex multiplication of $T$ are given by the roots of a sextic polynomial, which we have included in the Appendix. In light of Lemma 4.3, we conjecture that:
@@ 4.2 Extremisers of $\lambda_{n-1}$ with $\lambda_{n}$ repeated
[…]
@@
Thus if $n\geq 15$, we have $\lambda_{n-1}(M+E)\geq-\frac{n}{2\sqrt{3}}>-\frac{n-2}{3}$. But, the complements of $H_{a,b}$ have $\lambda_{n-1}(\overline{H_{a,b}})=-\frac{|V(\overline{H_{a,b}})|}{3}$, contradict the extremality of $\lambda_{n-1}(M+E)$. The authors have also verified for $5\leq n\leq 14$, there exists weighted graphs with penultimate eigenvalue lower than $\frac{n}{2\sqrt{3}}$.
@@
[…]
Let $\tilde{\mathbb{v}}_{n},\tilde{\mathbb{v}}_{n-1},\dots,\tilde{\mathbb{v}}_{1}$ be orthonormal eigenvectors of $M+E$, with corresponding eigenvalues $\tilde{\lambda}_{n}=\tilde{\lambda}_{n-1},\dots\tilde{\lambda}_{1}$. Then, consider the vector space $V=\braket{\mathbb{v,w}}\cap\braket{\mathbb{\tilde{v}}_{1},\dots\mathbb{\tilde{v}}_{n-1}}$. Note that $V$ has dimension at least 1, hence there exists some $\mathbb{z}\in V$ with $\mathbb{z}||=1$. Thus, if $\mathbb{x}=\sum_{i=1}^{n-1}a_{i}\mathbb{\tilde{v}}_{i}$, then:
@@
We remark that:
[…]
@@
[…]
which is less than or equal to 0 if and only if $S<0$ and $RS\geq Q^{2}$ (implying that $R\leq 0$). Identical logic applies when $R\neq 0$.
@@
[…]
@@
$E_{xy}=\begin{cases}-M_{xx}&x=y\\ 0&\text{otherwise.}\end{cases}$
@@
There exists a minimiser $M$ of $\lambda_{n-1}$ with at most 1 entry of each row and column not 0 nor $1$.
@@
Let $M$ be a minimiser of $\lambda_{n-1}$ with diagonal entries all zero and the least number of entries that are not 0 nor $1$. Suppose by way of contradiction some row (or equivalently column) of $M$ has at least 2 entries $M_{ab},M_{ac}\in(0,1)$ for some $a,b,c$ distinct. Consider symmetric matrices $E$ with 0 entries everywhere other than at $E_{ab},E_{ac},E_{bc}$ (and $E_{ba}$, $E_{ca}$, $E_{cb}$). Then, we have that:
@@
Regardless of the existence of a simple minimiser, there exists a symmetric $\{0,1\}$-matrix $N$ with zero entries along the diagonal such that for all $M\in\mathcal{S}_{n}$: